a. For classes of 162 students, find the mean and standard deviation for the number born on the 4th of July. Ignore leap years

b. For a class of 162 students, would two be an unusually high number who were born on the 4th of​ July?

a. The value of the mean is

μ=____.

​(Round to six decimal places as​ needed.)
Answer #1

a) P(Born at 4th July) = 1/365

Hence,

Mean = np = 162(1/365) = 0.443836

Standard deviation = /162(1/365)(364/365) = 0.665297

b) Maximum usual value = Mean + 2*SD = 0.443836 + 2*0.665297 = 1.77443

Since 2 is greater than the maximum usual value, it is an unusually high number.